OpenStudy (anonymous):

Integrate from 1 to 3: (2x^2+2)/sqrt(x^3+3x)

OpenStudy (campbell_st):

well I'd let \[u = x^3 + 3x\] so the boundary values become 4 to 36 then \[\frac{du}{dx} = 3x^2 + 3\] to get \[2x^2 + 2\] you need to let \[\frac{2}{3}du = (2x^2 + 2) dx\] so your problem is now \[\int\limits_{4}^{36} u^{-\frac{1}{2}} \times \frac{2}{3} du\] which can be written as \[\frac{2}{3} \int\limits_{4}^{36} u^{-\frac{1}{2}} du\] now integrate with respect to u and evaluate